James is 8% taller than John and wilson is 10% shorter than John. By what percentage is James taller than wilson?

Answer by fractalier(6550)

When you do problems like these, using 100 is good advice...
Thus let's call John's height 100...
Then James is 108 and Wilson is 90.
The percentage that James is taller than Wilson is given by
(108-90)/90 = 18/90 = .2 = 20%